The mayor of a community in MA has expressed opposition to having marijuana retail stores in the city. The public may or may not feel differently, and a local polling agency wants to find out. Surveying 580 residents, 310 are in support of allowing a marijuana retail outlet within the city, and 270 are opposed. Based on these data, determine if there is a clear and significant majority viewpoint.
Answer by Boreal(15235)

Ho: p=0.5
Ha: p NE 0.5
alpha=0.05 (will use that)prob.{reject Ho|Ho true}
phat=310/580=0.5345
z=(0.5345-0.500)/sqrt(0.5*0.5/580)
=0.0345/0.0207
=1.67
fail to reject Ho; insufficient evidence to say support is not 50% in the population.
p-value 0.096